r 16 (2.5-2.11) l and QUESTION 1 s connected to two oppositely charged parallel plates that are 10 cm apart. Calculate the electric force on a+200 microC point charge y 10 cm 00.1N 0.2N O 10 N 20 N 0.5 N QUESTION 2 17
Answer #1

V=100V ,d=10cm,

Relation between potential difference and electric field is

-dV/dr=E

E=100/10×10^-2=10^3N/C

F=qE=200×10^-6×10^3=0.2N